OpenStudy (anonymous):

what is the equation in standard form, of a parabola that models the values in the table. x|-2|0|4 F(x) |-7|3|-73

OpenStudy (kirbykirby):

You could approach this by either guessing values for the coefficients. Since you have "small" integer values to deal with, this is do-able if you have some good intuition. However, You can approach this more systematically: the equation of the parabola has the form \(f(x) = ax^2+bx+c\) , and they give you 3 x-values and 3-f(x) values. Since the parabola has 3 coefficients \(a, b, c\), you can find then by doing a system of 3 equations and 3 unknowns. However, you can reduce that to 2 equations and 2 unknowns very quicly by observing that for \(x=0\), \(f(0)=3\) which implies \(c=3\) ... this is because \(f(0)=a(0)^2+b(0)+c\\ 3 = 0 + 0 + c \implies c=3\) So, the rest can be found doing: \[-7=a(-2)^2+b(-2)+3\\ -73=a(4)^2+b(4)+3 \] which is a system of equations which you can solve for a and b